# zkFuzz: Foundation and Framework for Effective Fuzzing of Zero-Knowledge Circuits

### v1

@@ Abstract.
Our theoretical contribution is the Trace-Constraint Consistency Test (TCCT), a foundational language-independent formulation of ZK circuit bugs that defines bugs as discrepancies between the execution traces of the computation and the circuit constraints. TCCT captures both intermediate computations and program aborts, detecting bugs that elude prior tools.
@@
Our systems contribution is zkFuzz, a novel program mutation-based fuzzing framework for detecting TCCT violations. zkFuzz systematically mutates the computational logic of Zk programs guided by a novel fitness function, and injects carefully crafted inputs using tailored heuristics to expose bugs. We evaluated zkFuzz on 354 real-world ZK circuits written in Circom, a leading programming system for ZK development. zkFuzz successfully identified 66 bugs, including 38 zero-days—18 of which were confirmed by developers and 6 fixed, earning bug bounties.
Zero Knowledge Proof, Fuzzing, Circom
1. Introduction
@@
Zero-knowledge (ZK) circuits have emerged as a foundational technology for privacy-preserving computation with succinct proofs and efficient verification (Lavin et al., 2024; Morais et al., 2019; Hasan, 2019). Their adoption spans a variety of critical sectors, including anonymous cryptocurrencies (Ruj, 2024), confidential smart contracts (Yang and Li, 2020; Steffen et al., 2022), secure COVID-19 contact tracing protocols (Liu et al., 2020), privacy-preserving authentication (Baldimtsi et al., 2024), and verifiable machine learning systems (Xing et al., 2023; Zhang and Fan, 2024). Industry forecasts predict that the ZK market will reach \$10 billion by 2030, with Web3 applications alone expected to execute nearly 90 billion proofs (Protocol Labs, 2023).
@@
Manually constructing ZK circuits is complex and error-prone, so developers often use programming systems like Circom (Bellés-Muñoz et al., 2022), which allow them to express both the computation for deriving the secret witness from the input signals and the constraints of the finite-field arithmetic circuit (used to generate and verify proofs later) within one ZK program written in a high-level, domain-specific language (DSL). The DSL compiler then translates the program into (1) an executable (e.g., in WASM) for witness generation and (2) a set of constraints that define the circuit and support proof verification.
@@
Even with high-level programming systems, developing correct circuits remains notoriously difficult due to two core challenges. First, the computation and the constraints must align exactly, yet they operate in different domains: general-purpose logic vs. polynomial constraints optimized for efficiency. Developers must often manually translate logic into polynomial form, and any mismatch results in incorrect or unverifiable proofs (Wen et al., 2024). Second, the non-intuitive behavior of modular arithmetic in finite fields often confuses developers. For example, function iszero(x) should return $1$ if $x=0$ and 00 otherwise—but in a field of prime order $5$, expression iszero(2 + 3) evaluates to $1$, since $2+3\equiv 0\mod 5$. Such surprises often trip up developers unfamiliar with finite fields.
@@
Unsurprisingly, these challenges lead to vulnerabilities: under-constrained circuits that allow false proofs, and over-constrained circuits that reject valid witnesses. Both can result in serious security breaches (Chaliasos et al., 2024; Liang et al., 2025). For instance, an under-constrained bug in zkSync—a widely used zk-Rollup—allowed a malicious prover to extract \$1.9 billion worth of funds at the time of disclosure (Chainlight, 2023; Tang et al., 2024).
Existing methods for automatically detecting vulnerabilities in ZK programs face several limitations. Static analysis (Wen et al., 2024; of Bits, 2024) is primarily pattern-based and fails to capture deep semantic inconsistencies, suffering from high false positives, which waste developer time and lose their trust (Bessey et al., 2010). Formal methods (Isabel et al., 2024; Liu et al., 2024; Pailoor et al., 2023; Chen et al., 2024; Jiang et al., 2025) improve precision, but their reliance on SMT solvers limits scalability to real-world circuits with thousands of constraints. (Existing dynamic methods (Hochrainer et al., 2024; Leeb, 2024; Xiao et al., 2025; Fan et al., 2024) target compilation, proving, or verification systems, and do not extend to individual ZK programs.)
Furthermore, current formal definitions of ZK vulnerabilities are incomplete, leaving numerous bugs undetected. In particular, they ignore or mishandle intermediate computations (Pailoor et al., 2023; Chen et al., 2024; Jiang et al., 2025), missing an entire class of over-constrained bugs. They also fail to account for abnormal termination in computation—bugs that allow inputs to crash witness generation while still satisfying circuit constraints, despite such cases accounting for nearly 50% of real-world ZK bugs in our evaluation.
@@
Our first contribution in this paper is a new theoretical formulation, the Trace-Constraint Consistency Test (TCCT), that models the vulnerabilities in ZK programs as discrepancies between $(1)$ all possible execution traces that a computation may produce and $(2)$ the set of input, intermediate, and output values permitted by the circuit constraints. It captures both under-constrained, or soundness, vulnerabilities—when an invalid trace is allowed by the circuit constraints—and over-constrained, or completeness, vulnerabilities when a valid trace is incorrectly rejected. TCCT correctly accounts for intermediate computations, and is the first comprehensive model to define the semantics of over-constrained ZK vulnerabilities. It is also the first, to the best of our knowledge, to handle abnormal termination in witness computation. Our formulation is general and language-agnostic, laying a solid foundation for tools that detect both soundness and completeness ZK vulnerabilities.
@@
In each iteration, zkFuzz scores program mutants using a novel min-sum fitness function that estimates the likelihood of triggering a violation. For each input, it sums constraint errors across all constraints and assigns the mutant the minimum such sum across inputs. Unlike prior sum-only methods (Chen et al., 2022b), this approach is better suited for joint program and input fuzzing. zkFuzz then performs crossover, favoring lower-scoring mutants, and evaluates the resulting offspring on newly sampled inputs. This process repeats until a violation is found or a timeout occurs. Each iteration is independent once crossover mutants are generated, and zkFuzz further improves efficiency via heuristics that prioritize mutation of program statements and input regions most likely to cause bugs.
In our evaluation, we implement zkFuzz for Circom (Bellés-Muñoz et al., 2022), the most popular programming system for developing ZK circuits (Wen et al., 2024), and compare it with four state-of-the-art detection tools on 354 real-world ZK circuits. Our results show that zkFuzz consistently outperforms all other methods, finding 30% to 300% more bugs, without any false positives. In total, zkFuzz found 66 bugs, including 38 previously unknown zero-days—18 of which were confirmed by developers and 6 fixed. For example, 11 under-constrained bugs in zk-regex (zkemail, 2025), a popular ZK regex verification tool, were confirmed and awarded bug bounties. These bugs allow malicious provers to generate bogus proofs claiming that they possess a string with certain properties (e.g., a valide email address) without actually having such a string. Another confirmed bug in passport-zk-circuits (Rarimo, 2025), a ZK-based biometric passport project, allowed attackers to forge proofs of possessing data matching a required encoding. Our code is publicly available at https://github.com/Koukyosyumei/zkFuzz.
In summary, this paper makes the following key contributions:
  • Introduce TCCT, a language-independent formal framework that captures both under- and over-constrained ZK bugs through trace-constraint inconsistencies.
  • Develop zkFuzz, a dynamic analysis tool that detects TCCT violations by jointly mutating programs and inputs.
@@
  • Evaluate zkFuzz on 354 real-world ZK circuits, finding 30–300% more bugs than prior tools with zero false positives.
Figure 1. Overview of ZK proof systems. The circuit compiler processes a ZK program into a witness computation program and circuit constraints. Prover executes the witness program to obtain a trace (witness and public values) and creates a proof using the proving system. The verifier validates the proof using the verification system with the public output, the constraint, and the proof. Our fuzzer checks the inconsistencies between the computation logic and the circuit constraints in the ZK program.
2. Background
This section provides an overview of ZK Proof systems (§ 2.1), vulnerabilities in ZK circuits (§ 2.2), and Circom, the most popular programming system for ZK circuits (Belles-Munoz et al., 2023; Wen et al., 2024) (§ 2.3).
2.1. ZK Proof Systems
ZK proofs enable a prover to convince a verifier of the validity of a statement (e.g., knowledge of a secret value) without revealing any information beyond the statement’s truth (Li and McMillin, 2014). Formally, ZK proof systems satisfy three properties:
@@ 2.1. ZK Proof Systems
These properties enable the construction of privacy-preserving verifiable systems, such as anonymous cryptocurrencies (e.g., Zcash’s zk-SNARKs), confidential smart contracts, and verifiable voting protocols. Modern applications extend to blockchain scaling (e.g., ZK-Rollups) and verification of machine-learning integrity (Lavin et al., 2024; Garg et al., 2023; Abbaszadeh et al., 2024).
At a high level, a ZK system operates on arithmetic circuits that perform computations over a finite field, where all variables and operations (e.g., addition, multiplication) are defined modulo a large prime. It exposes two primitives: $\mathrm{Prove}$ and $\mathrm{Verify}$. The ZK proof process begins with the prover running $\mathrm{Prove}$ to generate a proof: $\pi\leftarrow\mathrm{Prove}(\mathcal{C},w)$ where $\mathcal{C}$ is the circuit and $w$ the prover’s secret value, known as the witness. Importantly, the proof $\pi$ is constructed to reveal no information about $w$. Once $\pi$ is generated, the prover sends it to the verifier who then calls $\mathrm{Verify}(\mathcal{C},\pi)$ which returns either $accept$ or $reject$. For performance reasons, it is often desirable for the proof to be succinct and non-interactive.
@@
Manually constructing circuits for such computations and determining which values belong in the witness is complex and error-prone, so developers often use ZK programming systems (Lavin et al., 2024) like Circom, which allow them to express the witness computation and the circuit constraints in a single ZK program written in a high-level DSL. The DSL compiler then translates the program into a computation program $\mathcal{P}$ and the corresponding set of circuit constraints $\mathcal{C}$. When $\mathcal{P}$ executes, it produces a full execution trace, including not only the input $x$ and output $y$, but also important intermediate values $z$, to match the required constraint format. For example, computing $y=x^{4}$ typically requires two quadratic constraints: $z=x^{2}$ and $y=z^{2}$. The resulting execution trace $(x,z,y)$ becomes the witness (typically, $y$ is public, while $x$ and $z$ are secret; however, developers sometimes make $x$ public as well, depending on the application and desired level of privacy.) Fig. 1 illustrates this process, along with the subsequent use of proving and verification systems like snarkjs (iden3, 2018).
Although some systems like Noir (Aztec Network, 2024) support automatic generation of constraints from the computations, the resulting constraints are often significantly less efficient, increasing the number of constraints by 3x to 300x (noir lang, 2025). In the remainder of our paper, we focus on practical ZK systems that produce efficient constraints.
2.2. Overview of Vulnerabilities in ZK Programs
Even with high-level programming systems, general computations often involve non-quadratic operations that cannot be directly mapped to the required constraint format. As a result, it is the developer’s responsibility to ensure that the intended computation $\mathcal{P}$ and the corresponding circuit constraints $\mathcal{C}$ are aligned; any mismatch may lead to incorrect or unverifiable proofs. Unfamiliarity with finite-field arithmetic further complicates matters, as its modular behavior introduces unexpected corner cases. Consequently, ZK programs are notoriously difficult to get right (Pailoor et al., 2023; Chen et al., 2024) and may contain the following two classes of vulnerabilities.
Under-Constrained Circuits. $\mathcal{C}$ is too loose for $\mathcal{P}$, allowing malicious provers to convince verifiers that ”I know the input whose corresponding output is $y$,” even when they do not. Such vulnerabilities violate the soundness property.
@@ 2.2. Overview of Vulnerabilities in ZK Programs
Code 1 and 2 show the computation and constraints, respectively, for an under-constrained vulnerability caught by zkFuzz in a real-world 1-bit right-shift program from the circom-monolith library (Shukla, 2023). Since right shift is not a quadratic operation and cannot be directly encoded as a constraint, the developer introduces an intermediate value b, defined as the difference between x and 2y, where y = x >> 1. Unfortunately, due to quirks in finite-field arithmetic, multiple assignments to (y, b) can satisfy the constraints for a given input x. For example, in the finite field $\mathbb{F}_{11}$, if x = 7, the expected output is y = 3 and b = 1. Yet, y = 9, b = 0 also satisfy the constraints, because $7-9\cdot 2=-11\equiv 0\bmod 11$. The circuit is thus under-constrained, allowing an invalid proof.
@@
The safe implementation of 1-bit right shift using the bit array can be found in Appendix B.
@@
Out-of-Scope Vulnerabilities. zkFuzz focuses on soundness and completeness vulnerabilities in ZK proofs that arise from developer mistakes in tying constraints to computation. The following types of vulnerabilities are out of scope: (1) vulnerabilities in the underlying compilation, proving, and verification systems (e.g., the Frozen Heart vulnerability caused by an insecure implementation of the Fiat-Shamir transformation (Tang et al., 2024)); (2) bugs solely in the computation logic, such as unused inputs in hashing (Wen et al., 2024) (such unused inputs may be acceptable depending on the use case, as in computing the trace of a matrix where only diagonal elements are used); and (3) bugs solely in ZK protocol design that leak information and violate the zero-knowledge property (Chaliasos et al., 2024), inline with all prior detectors (of Bits, 2024; Pailoor et al., 2023; Jiang et al., 2025; Chen et al., 2024; Wen et al., 2024)—what constitutes a secret is highly application-dependent: e.g., in Validity Rollups (Chen et al., 2022a), ZK proofs are used to accelerate transactions rather than to conceal them.
2.3. Circom Primer

### v2

@@ Abstract
Our theoretical contribution is the Trace-Constraint Consistency Test (TCCT), a foundational, language-independent formulation of ZK circuit bugs. TCCT provides a unified semantics that subsumes prior definitions and captures both under- and over-constrained vulnerabilities, exposing the full space of ZK bugs that elude prior tools.
@@
Our systems contribution is zkFuzz, a novel program mutation-based fuzzing framework for detecting TCCT violations. zkFuzz systematically mutates the computational logic of Zk programs guided by a novel fitness function, and injects carefully crafted inputs using tailored heuristics to expose bugs. We evaluated zkFuzz on 452 real-world ZK circuits written in Circom, a leading programming system for ZK development. zkFuzz successfully identified 85 bugs, including 59 zero-days—39 of which were confirmed by developers and 14 fixed, including bugs undetectable by prior works due to their fundamentally limited formulations, earning thousands of bug bounties. Our preliminary research on Noir, another emerging DSL for ZK circuit, also demonstrates the feasibility of zkFuzz to support multiple DSLs.
11footnotetext: This paper has been accepted at IEEE Symposium on Security and Privacy (S&P) 2026.
1 Introduction
@@
Zero-knowledge (ZK) circuits have emerged as a foundational technology for privacy-preserving computation with succinct proofs and efficient verification [1, 2, 3]. Their adoption spans a variety of critical sectors, including anonymous cryptocurrencies [4], confidential smart contracts [5, 6], secure COVID-19 contact tracing protocols [7], privacy-preserving authentication [8], and verifiable machine learning systems [9, 10]. Industry forecasts predict that the ZK market will reach \$10B by 2030, with Web3 applications alone expected to execute nearly 90 billion proofs [11].
@@
Manually constructing ZK circuits is complex and error-prone, so developers often use programming systems like Circom [12], which allow them to express both the computation for deriving the secret witness from the input signals and the constraints of the finite-field arithmetic circuit (used to generate and verify proofs later) within one ZK program written in a high-level, domain-specific language (DSL). The DSL compiler then translates the program into (1) an executable (e.g., in WASM) for witness generation and (2) a set of constraints that define the circuit and support proof verification.
@@
Even with high-level programming systems, developing correct circuits remains notoriously difficult due to two core challenges. First, the computation and the constraints must align exactly, yet they operate in different domains: general-purpose logic vs. polynomial constraints optimized for efficiency. Developers must often manually translate logic into polynomial form, and any mismatch results in incorrect or unverifiable proofs [13]. Second, the non-intuitive behavior of modular arithmetic in finite fields often confuses unfamiliar developers.
@@
Unsurprisingly, these challenges lead to vulnerabilities: under-constrained circuits that allow false proofs, and over-constrained circuits that reject valid witnesses. Both can result in serious security breaches [14, 15]. For instance, an under-constrained bug in zkSync—a widely used zk-Rollup—allowed a malicious prover to extract \$1.9 billion worth of funds at the time of disclosure [16, 17]. While the risk of under-constrained circuits is widely acknowledged [18, 15, 14], over-constrained circuits can also cause severe denial-of-service vulnerabilities for honest provers. For example, one over-constrained bug was recently discovered in RISC-Zero, the most widely used zkVM [19], which silently rejected valid execution traces, effectively compromising the integrity of the virtual machine.
Existing methods for detecting vulnerabilities in ZK programs are limited. Static analysis tools rely on patterns and miss deep semantic issues, leading to high false positives that erode developer trust [13, 20, 21]. Formal methods offer better precision but don’t scale to large circuits due to SMT solver bottlenecks [22, 23, 18, 24, 25]. Dynamic approaches focus on compilation or verification systems and don’t generalize to individual ZK programs [26, 27, 28, 29].
Furthermore, existing formal definitions of ZK vulnerabilities are incomplete, leaving many bugs undetected. In particular, prior work focused only on constraints, ignoring program semantics, and defined bugs solely as non-determinism in constraints—situations where multiple outputs for the same input all satisfy the constraints [18, 24, 25]. This omission misses two major classes: over-constrained bugs and under-constrained bugs due to abnormal termination, where inputs can crash witness generation while still satisfying circuit constraints. Together, these overlooked categories account for nearly 50% of real-world ZK bugs in our evaluation.
@@
Our first contribution in this paper is a new theoretical formulation, the Trace-Constraint Consistency Test (TCCT), that models the vulnerabilities in ZK programs as discrepancies between $(1)$ all possible execution traces that a computation may produce and $(2)$ the set of input, intermediate, and output values permitted by the circuit constraints. TCCT captures both under-constrained, or soundness, vulnerabilities—when an invalid trace is accepted by the circuit constraints—and over-constrained, or completeness, vulnerabilities—when a valid trace is incorrectly rejected. TCCT provides two major advantages over prior definitions. First, it offers a unified semantics: TCCT is the first framework to capture both under- and over-constrained vulnerabilities within the same formal model, correctly accounting for intermediate computations and handling abnormal termination in witness generation. This unification subsumes prior definitions as special cases and exposes the full space of ZK vulnerabilities. Second, TCCT is language-agnostic: it operates at the level of trace/constraint properties rather than DSL syntax. While our implementation, zkFuzz, currently supports Circom and has very preliminary support for Noir, many DSLs already compile to standardized formats (e.g., R1CS), and by similarly defining a universal format for execution traces, zkFuzz can readily extend across diverse ZK languages.
@@
In each iteration, zkFuzz scores program mutants using a novel min-sum fitness function that estimates the likelihood of triggering a violation. For each input, it sums constraint errors across all constraints and assigns the mutant the minimum such sum across inputs. Unlike prior sum-only methods [30], this approach is better suited for joint program and input fuzzing. zkFuzz then performs crossover, favoring lower-scoring mutants, and evaluates the resulting offspring on newly sampled inputs. This process repeats until a violation is found or a timeout occurs. Each iteration is independent once crossover mutants are generated, and zkFuzz further improves efficiency via heuristics that prioritize mutation of program statements and input regions most likely to cause bugs.
In our evaluation, we implement zkFuzz for Circom [12], the most popular programming system for developing ZK circuits [13], and compare it with four state-of-the-art detection tools on 452 real-world ZK circuits. Our results show that zkFuzz consistently outperforms all other methods, finding 30% to 300% more bugs, without any false positives. In total, zkFuzz found 85 bugs, including 59 previously unknown zero-days—39 of which were confirmed by developers and 14 fixed. For example, 11 under-constrained bugs in zk-regex [31], a popular ZK regex verification tool, were confirmed and awarded bug bounties. These bugs allow malicious provers to generate bogus proofs claiming that they possess a string with certain properties (e.g., a valid email address) without actually having such a string. Another confirmed bug in passport-zk-circuits [32], a ZK-based biometric passport project, allowed attackers to forge proofs of possessing data matching a required encoding. We also prototype zkFuzz for Noir [33], another emerging DSL, and evaluate it on real Noir circuits. Our code is publicly available at https://github.com/Koukyosyumei/zkFuzz.
In summary, we make the following key contributions:
  • Introduce TCCT, a language-independent formal framework capturing both under- and over-constrained ZK bugs through trace-constraint inconsistencies.
  • Develop zkFuzz, a fuzzer that detects TCCT violations by jointly mutating programs and inputs.
@@
  • Evaluate zkFuzz on 452 real-world ZK circuits written in Circom, finding 30–300% more bugs than prior tools with zero false positives.
  • Implement a preliminary prototype for Noir, another popular DSL for ZK circuits not supported by existing tools, and demonstrate that zkFuzz can detect a real ZK bug reported previously in Noir circuits.
Figure 1: Overview of ZK proof systems. The circuit compiler processes a ZK program into a witness computation program and circuit constraints. Prover executes the witness program to obtain a trace (witness and public values) and creates a proof using the proving system. The verifier validates the proof using the verification system with the public output, the constraint, and the proof. Our fuzzer checks the inconsistencies between the computation logic and the circuit constraints in the ZK program.
2 Background
This section provides an overview of ZK Proof systems (§ 2.1), bugs in ZK circuits (§ 2.2), and Circom, the most popular programming system for ZK circuits [13] (§ 2.3).
2.1 ZK Proof Systems
ZK proofs enable a prover to convince a verifier of the validity of a statement without revealing any information beyond the statement’s truth [34]. Formally, ZK proof systems satisfy three properties:
@@ 2.1 ZK Proof Systems
These properties enable the construction of privacy-preserving verifiable systems, such as anonymous cryptocurrencies (e.g., Zcash’s zk-SNARKs), confidential smart contracts, and verifiable voting protocols. Modern applications extend to blockchain scaling (e.g., ZK-Rollups) and verification of machine-learning integrity [1, 35, 36].
At a high level, a ZK system operates on arithmetic circuits that perform computations over a finite field, where all variables and operations (e.g., addition, multiplication) are defined modulo a large prime. It exposes two primitives: $\mathrm{Prove}$ and $\mathrm{Verify}$. The ZK proof process begins with the prover running $\mathrm{Prove}$ to generate a proof: $\pi\leftarrow\mathrm{Prove}(\mathcal{C},w)$ where $\mathcal{C}$ is the circuit and $w$ the prover’s secret value, known as the witness. Importantly, the proof $\pi$ is constructed to reveal no information about $w$. Once $\pi$ is generated, the prover sends it to the verifier who then calls $\mathrm{Verify}(\mathcal{C},\pi)$ which returns either acceptaccept or rejectreject. For performance reasons, it is often desirable for the proof to be succinct and non-interactive.
@@
Manually constructing circuits for such computations and determining which values belong in the witness is complex and error-prone, so developers often use ZK programming systems [1] like Circom, which allow them to express the witness computation and the circuit constraints in a single ZK program written in a high-level DSL. The DSL compiler then translates the program into a computation program $\mathcal{P}$ and the corresponding set of circuit constraints $\mathcal{C}$. When $\mathcal{P}$ executes, it produces a full execution trace, including not only the input $x$ and output $y$, but also important intermediate values $z$, to match the required constraint format. For example, computing $y=x^{4}$ typically requires two quadratic constraints: $z=x^{2}$ and $y=z^{2}$. The resulting execution trace $(x,z,y)$ becomes the witness (typically, $y$ is public, while $x$ and $z$ are secret; however, developers sometimes make $x$ public as well, depending on the application and desired level of privacy.) Fig. 1 illustrates this process, along with the subsequent use of proving and verification systems like snarkjs [37].
Although some systems like Noir [33] support automatic generation of constraints from the computations, the resulting constraints are often significantly less efficient, increasing the number of constraints by 3x to 300x [38]. Consequently, many Noir projects use unsafe mode, where programmers manually write constraints like Circom, introducing the risk of discrepancy between the computation program and constraints. In the remainder of our paper, we focus on practical ZK systems that produce efficient constraints.
2.2 Overview of Vulnerabilities in ZK Programs
Even with high-level programming systems, general computations often involve non-quadratic operations that cannot be directly mapped to the required constraint format. As a result, it is the developer’s responsibility to ensure that the intended computation $\mathcal{P}$ and the corresponding circuit constraints $\mathcal{C}$ are aligned; any mismatch may lead to incorrect or unverifiable proofs. Unfamiliarity with finite-field arithmetic further complicates matters, as its modular behavior introduces unexpected corner cases. Consequently, ZK programs are notoriously difficult to get right [18, 24] and may contain the following two classes of vulnerabilities.
  • Under-Constrained Circuits. $\mathcal{C}$ is too loose for $\mathcal{P}$, allowing malicious provers to convince verifiers that ”I know the input whose corresponding output is yy,” even when they do not. Such vulnerabilities violate the soundness property.
@@ 2.2 Overview of Vulnerabilities in ZK Programs
Code 1 and 2 show the computation and constraints, respectively, for an under-constrained vulnerability caught by zkFuzz in a real-world 1-bit right-shift program from the circom-monolith library [39]. Since right shift is not a quadratic operation and cannot be directly encoded as a constraint, the developer introduces an intermediate value b, defined as the difference between x and 2y, where y = x >> 1.
Unfortunately, due to quirks in finite-field arithmetic, multiple assignments to (y, b) can satisfy the constraints for a given input x. For example, in the finite field $\mathbb{F}_{11}$, if x = 7, the expected output is y = 3 and b = 1. Yet, y = 9, b = 0 also satisfy the constraints, because $7-9\cdot 2=-11\equiv 0\bmod 11$. The circuit is thus under-constrained, allowing an invalid proof.
@@
Out-of-Scope Vulnerabilities. zkFuzz focuses on soundness and completeness vulnerabilities in ZK proofs that arise from developer mistakes in tying constraints to computation. The following types of vulnerabilities are out of scope: (1) vulnerabilities in the underlying compilation, proving, and verification systems (e.g., the Frozen Heart vulnerability caused by an insecure implementation of the Fiat-Shamir transformation [17]); (2) bugs solely in the computation logic, such as unused inputs in hashing [13] (such unused inputs may be acceptable depending on the use case, as in computing the trace of a matrix where only diagonal elements are used); and (3) bugs solely in ZK protocol design that leak information and violate the zero-knowledge property [14], inline with all prior detectors [20, 18, 25, 24, 13]—what constitutes a secret is highly application-dependent: e.g., in Validity Rollups [40], ZK proofs are used to accelerate transactions rather than to conceal them.
2.3 Circom Primer